OpenStudy (anonymous):

if the derivative of y=1+3x^2/3 at x=-8 is y'= 2x^ -1/3 How do we get the x-value appropriate rule 2(-8)^-1/3 how do i simply from there please provide steps and rules thanks.

OpenStudy (anonymous):

are you asking how you would evaluate 2(-8)^-1/3

OpenStudy (anonymous):

Yes on my book it shows answer to be -1. I am lost as to how was that simplied.

OpenStudy (anonymous):

the 3rd root of 8 is 2

OpenStudy (aum):

\[ \large 2(-8)^{-1/3} = \frac{2}{\sqrt[3]{-8}} = ? \] What is the cube root of -8?

OpenStudy (aum):

-8 = (-2)^3 So cube root of -8 is -2. 2 / (-2) = -1.

OpenStudy (anonymous):

aum thanks your so awesome. All easy steps and i just cant figure this algebra out. THANKS AGAIN!

OpenStudy (aum):

You are welcome.

OpenStudy (aum):

\[ \large 2(-8)^{-1/3} = \frac{2}{(-8)^{1/3}} = \frac{2}{\{(-2)^3\}^{1/3}} = \frac{2}{(-2)^{3*1/3}} = \\ \large \frac{2}{(-2)^{1}} = \frac{2}{-2} = -1 \]